# Kerr-MOG-(A)dS black hole and its shadow in scalar-tensor-vector gravity theory

### v2

@@ V CONCLUSIONS AND OUTLOOKS
  • Size: As the cosmological constant increases, the size of the black hole shadow decreases. But the size of the shadow does not appear to be monotonic as $\alpha$ changes. For example, it would reach a maximum at $\alpha\approx 1.815$ with $\Lambda=10^{-3}$.
@@
There are two promising further topics to be pursued in the future. An intriguing topic is to explore other observable measurements of the shadow of the Kerr-MOG-dS black hole. To prevent certain shadow degeneracy caused by black hole parameters, we can explore other distortion parameters, such as the oblateness parameter $K_{s}$, and the ”thickness” parameter $T$ of the shadow Wang et al. (2018b). In order to visualize the shadow in a real astronomical setting, one should rely on the emissions originating from the accretion disk surrounding the black hole Zhang et al. (2021, 2022). And the other is to investigate the properties of the Kerr-MOG-AdS black hole, such as black hole thermodynamics Christodoulou (1970); Bardeen et al. (1973); Wu and Wu (2019); Cong et al. (2021); Ahmed et al. (2023a), thermodynamic topological classification Wei et al. (2022); Wu (2023a); Wu and Wu (2023); Wu (2023b), null hypersurface caustics Al Balushi and Mann (2019); Imseis et al. (2021); Wu et al. (2021), phase transition criticality Kubiznak and Mann (2012); Ahmed et al. (2023b); Wu et al. (2024), Joule-Thomson expansion Ökcü and Aydıner (2018), etc.
@@ Acknowledgements.
The authors sincerely thank Professor Songbai Chen for his contributions to this subject and Dr. Long Fen for the profound discussions. This work was supported by the National Natural Science Foundation of China under Grants No. 12122504, No. 12375046, No. 12035005, No. 12205243, No. 12375053, by the innovative research group of Hunan Province under Grant No. 2024JJ1006, by the Natural Science Foundation of Hunan Province under grant No. 2023JJ30384, by the Hunan provincial major sci-tech program under grant No.2023zk1010, by the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347, and by the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028.
@@ Appendix A Energy-momentum tensor
ϕ¯=¯italic-$\bar{\phi}=$ $-\frac{\sqrt{G_{\text{N}}\alpha}M\bar{r}}{\Sigma}\left(\frac{\sin^{2}\bar{\vartheta}}{\chi}+\cos^{2}\bar{\vartheta}\right)d\bar{t}-\frac{h\sqrt{G_{\text{N}}\alpha}M\bar{r}}{\Delta_{r}\Delta_{\bar{r}}}d\bar{r}+\frac{\sqrt{G_{\text{N}}\alpha}M\bar{r}}{\chi\Sigma}a\sin^{2}\bar{\vartheta}d\bar{\varphi},$ (56)
@@
$T_{00}=$ $-\frac{1}{24\pi\chi^{2}\Sigma^{3}}G_{\text{N}}M^{2}\alpha\Delta_{\bar{\vartheta}}\left[(3-\bar{r}^{2}\Lambda)\Sigma_{2}+3h\Delta_{\bar{\vartheta}}\right],$ $T_{01}=-\frac{1}{8\pi\chi\Delta_{\bar{r}}\Sigma^{2}}G_{\text{N}}M^{2}\alpha h\Delta_{\bar{\vartheta}},$ (57)
$T_{11}=$ $-\frac{1}{8\pi\Delta_{\bar{r}}^{2}\Sigma}G_{\text{N}}M^{2}\alpha\left(\Delta_{r}-2\Delta_{\bar{r}}\right),$ $T_{03}=\frac{1}{8\pi\chi^{2}\Sigma^{3}}G_{\text{N}}M^{2}\alpha\Delta_{\bar{\vartheta}}\left(\Delta_{r}+\Delta_{\bar{r}}\right)a\sin^{2}\bar{{\vartheta}},$
$T_{22}=$ $-\frac{1}{8\pi\Delta_{\bar{\vartheta}}\Sigma}G_{\text{N}}M^{2}\alpha,$ $T_{13}=\frac{1}{8\pi\chi\Delta_{\bar{r}}\Sigma^{2}}G_{\text{N}}M^{2}\alpha ha\sin^{2}\bar{\vartheta},$

### v3

@@ V CONCLUSIONS AND OUTLOOKS
  • Size: As the cosmological constant increases, the size of the black hole shadow decreases. But the size of the shadow does not appear to be monotonic as $\alpha$ changes. For example, it would reach a maximum at $\alpha\approx 1.815$ with $\tilde{\Lambda}=10^{-3}$.
@@
There are two promising further topics to be pursued in the future. An intriguing topic is to explore other observable measurements of the shadow of the Kerr-MOG-dS black hole. To prevent certain shadow degeneracy caused by black hole parameters, we can explore other distortion parameters, such as the oblateness parameter $K_{s}$, and the ”thickness” parameter $T$ of the shadow Wang et al. (2018b). In order to visualize the shadow in a real astronomical setting, one should rely on the emissions originating from the accretion disk surrounding the black hole Zhang et al. (2021, 2022). And the other is to investigate the properties of the Kerr-MOG-AdS black hole, such as black hole thermodynamics Christodoulou (1970); Bardeen et al. (1973); Wu and Wu (2019); Cong et al. (2021); Ahmed et al. (2023a); Liu et al. (2024d), thermodynamic topological classification Wei et al. (2022); Wu (2023a); Wu and Wu (2023); Wu (2023b), null hypersurface caustics Al Balushi and Mann (2019); Imseis et al. (2021); Wu et al. (2021), phase transition criticality Kubiznak and Mann (2012); Ahmed et al. (2023b); Wu et al. (2024), Joule-Thomson expansion Ökcü and Aydıner (2018), etc.
@@ Acknowledgements.
The authors sincerely thank Professor Songbai Chen for his contributions to this subject and Dr. Fen Long for the profound discussions. This work was supported by the National Natural Science Foundation of China under Grants No. 12122504, No. 12375046, No. 12035005, No. 12205243, No. 12375053, by the innovative research group of Hunan Province under Grant No. 2024JJ1006, by the Natural Science Foundation of Hunan Province under grant No. 2023JJ30384, by the Hunan provincial major sci-tech program under grant No.2023zk1010, by the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347, and by the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028.
@@ Appendix A Energy-Momentum Tensor
ϕ¯=¯italic-$\bar{\phi}=$ $-\frac{\sqrt{G_{\text{N}}\alpha}M\bar{r}}{\Sigma}\left(\frac{\sin^{2}\bar{\vartheta}}{\chi}+\cos^{2}\bar{\vartheta}\right)d\bar{t}$ (56)
$-\frac{h\sqrt{G_{\text{N}}\alpha}M\bar{r}}{\Delta_{r}\Delta_{\bar{r}}}d\bar{r}+\frac{\sqrt{G_{\text{N}}\alpha}M\bar{r}}{\chi\Sigma}a\sin^{2}\bar{\vartheta}d\bar{\varphi},$
@@
$T_{00}=$ $-\frac{1}{24\pi\chi^{2}\Sigma^{3}}G_{\text{N}}M^{2}\alpha\Delta_{\bar{\vartheta}}\left[(3-\bar{r}^{2}\Lambda)\Sigma_{2}+3h\Delta_{\bar{\vartheta}}\right],$ (57)
$T_{01}=$ $-\frac{1}{8\pi\chi\Delta_{\bar{r}}\Sigma^{2}}G_{\text{N}}M^{2}\alpha h\Delta_{\bar{\vartheta}},$
$T_{11}=$ $-\frac{1}{8\pi\Delta_{\bar{r}}^{2}\Sigma}G_{\text{N}}M^{2}\alpha\left(\Delta_{r}-2\Delta_{\bar{r}}\right),$
$T_{03}=$ $\frac{1}{8\pi\chi^{2}\Sigma^{3}}G_{\text{N}}M^{2}\alpha\Delta_{\bar{\vartheta}}\left(\Delta_{r}+\Delta_{\bar{r}}\right)a\sin^{2}\bar{{\vartheta}},$
$T_{22}=$ $-\frac{1}{8\pi\Delta_{\bar{\vartheta}}\Sigma}G_{\text{N}}M^{2}\alpha,$
$T_{13}=$ $\frac{1}{8\pi\chi\Delta_{\bar{r}}\Sigma^{2}}G_{\text{N}}M^{2}\alpha ha\sin^{2}\bar{\vartheta},$
@@
Appendix B Ray-Tracing Method
In this appendix, we aim to provide a concise and useful overview of the numerical ray-tracing method used in our manuscript.
For Kerr-MOG-dS spacetime, the 4-momentum vector of a photon is written as
$p^{a}=(\dot{t},\dot{r},\dot{\vartheta},\dot{\varphi}),$ (58)
where the dot denotes the derivative with respect to the affine parameter $\tau$. Given that $g_{ab}p^{a}p^{b}=0$, we can define a dual vector
$q_{a}=g_{ab}p^{b},$ (59)
which satisfies $q_{a}p^{a}=0$. The geodesic equations, in the Hamiltonian canonical formalism, are given by
$\dot{q}_{a}=-\frac{\partial H}{\partial x^{a}},$ $\dot{x}^{a}=\frac{\partial H}{\partial q_{a}}$ (60)
where $H$ is the Hamiltonian of the null particle and $x^{a}$ are the 4-positions of the geodesic. These are first-order differential equations with respect to the canonical variables $(x^{a},q_{a})$, making them very convenient for numerical geodesic evolution.
Just like the approach in Sec. IV.1, we assume that an observer is located at ${r_{0},\vartheta_{0}}$ in the coordinates $\{t,r,\vartheta,\varphi\}$ and has chosen an orthonormal tetrad the same as in equation (41). In the observer’s frame, the tangent vector of the null geodesic can be expressed as we can rewrite equation (43) in this form,
$\dot{\lambda}=|\overrightarrow{OP}|\gamma(\sin\theta\cos\psi e_{1}+\sin\theta\sin\psi e_{2}+\cos\theta e_{3}-e_{0}),$
where $|\overrightarrow{OP}|$ denotes the tangent vector of the null geodesic at the point $O$ in the three-dimensional space. Then, we can re-obtain the stereographic projection from the celestial sphere onto a plane
$x_{P^{\prime}}=$ $-2|\overrightarrow{OP}|\tan\frac{\theta}{2}\sin\psi,$ (61)
$y_{P^{\prime}}=$ $-2|\overrightarrow{OP}|\tan\frac{\theta}{2}\cos\psi,$
$P^{\prime}$ represents the projection of point $P$ in the celestial coordinate system onto the Cartesian coordinate system. We can set up standard Cartesian coordinates with the origin $O^{\prime}$ , as shown in Fig. 9.
Figure 9: Illustration of the pixels. A standard Cartesian coordinate with the origin $O^{\prime}$ is set up.
Next, let us discuss the field of view. It is characterized by the angles in the planes $yO^{\prime}z$ and $xO^{\prime}z$. For convenience, we consider these two angles are equal. An example of the field of view is shown in Fig. 10.
Figure 10: The field of view.
To determine the dimensions of the square screen, we need to calculate its length. The length $L$ can be expressed as:
$L=2|\overrightarrow{OP}|\tan\frac{\alpha_{\text{~{}fov}}}{2}.$ (62)
We consider $n\times n$ pixels on the screen, where the length occupied per pixel is given by
$\ell=\frac{2|\overrightarrow{OP}|}{n}\tan\frac{\alpha_{\text{~{}fov}}}{2}.$ (63)
The pixels are labeled by $(i,j)$, with the pixel at the bottom left corner being $(1,1)$ and the pixel at the top right corner being $(n,n)$. The indices $i$ and $j$ range from $1$ to $n$. For the Cartesian coordinates of the center point of a pixel, we have
$x_{P^{\prime}}=\ell\left(i-\frac{n+1}{2}\right),$ $y_{P^{\prime}}=\ell\left(j-\frac{n+1}{2}\right).$ (64)
By comparing equations (61) and (64), we can establish a relationship between $(i,j)$ and $(\theta,\Psi)$ as follows:
$\tan\Psi$ $=\frac{j-(n+1)/2}{i-(n+1)/2},$ (65)
$\tan\frac{\theta}{2}$ $=\tan\frac{\alpha_{\text{~{}fov}}}{2}\frac{\sqrt{[i-(n+1)/2]^{2}+[j-(n+1)/2]^{2}}}{n}.$ (66)
Figure 11: Illustration of our spherical light source at infinity.
Finally, we state that we use an extended source to illuminate the system. The model, illustrated in Fig. 11, displays a cutaway view of the photosphere, exposing its internal structure. We segment the sphere into a grid network formed by lines of latitude and longitude, with an interval of $\pi/18$ between each pair of adjacent lines. To render an image of the black hole, we assign a specific color to each grid segment. The extended light source is divided into four distinct colors. The color of each pixel on the image is determined by tracing the path of photons from the corresponding points on the extended source. Areas that appear dark signify that the photons have been absorbed by the black hole.